OpenStudy (anonymous):

Find the next three terms in each pattern: 100, 81, 64, 49, 36, ... -3/4, 1/4, -1/6, 1/9, ...

OpenStudy (mathmath333):

\(100, 81, 64, 49, 36, ...\\ \implies 10^2,9^2,8^2,7^2,6^2........\)

OpenStudy (anonymous):

^

OpenStudy (mathmath333):

second one is tricky ,lol

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mathmath333):

it might be this but m not sure \(\Large\tt -\dfrac{2^{-3}}{3^{-1}},\dfrac{2^{-2}}{3^0},-\dfrac{2^{-1}}{3^1},\dfrac{2^0}{3^2}..........\)